A ball rolls off the edge of a horizontal table top 2.3 ft high. It strike the floor at a point 5.11 ft horizontally away from the edge of the table  a) For how long was the ball in air  b) what was the speed at the instant it left the table.


1
Expert's answer
2021-05-05T08:55:48-0400

a.) We can use,


"h = \dfrac{1}{2}gt^2"


Before that converting heights into metres

"h_{AB} = 2.3 ft = 0.70m"

"h_{BC} = 1.55m"


"h_{AB} = \dfrac{1}{2}g(t_{AC})^2"


"t_{AC} = \sqrt{\dfrac{2h_{AB}}{g}}"


"= \sqrt{\dfrac{2 \times0.70}{9.8}}"


"= 0.37 s"


b.)The speed at the instant it left the table.


"v = \dfrac{BC}{t_{AC}} = \dfrac{1.55}{0.37} = 4.18 m/s"


The diagram of this situation is :